OpenStudy (anonymous):

(b+1over3)(b+1 over 5)

OpenStudy (anonymous):

\[(b+\frac{1}{3})(b+\frac{1}{5})\] \[b^2+\frac{1}{5}b+\frac{1}{3}b+\frac{1}{3}\times \frac{1}{5}\] \[b^2+\frac{8}{15}b+\frac{1}{15}\]